OpenStudy (anonymous):

xy = -12 What is the constant of variation for this relationship? a. -1.2 b. -12 c. -1/4 d. -0.12

Nnesha (nnesha):

equation for direct variation is y = kx and for inverse variation is = y = k/x now in the equation what variable r u missing

OpenStudy (anonymous):

x

Nnesha (nnesha):

 both equation have 3 variable x y and k x and y Valeria are on left side

Nnesha (nnesha):

\[\large\rm y =\frac{ k }{ x }\] if you solve this for k you should have x y = k now what is k in that equation ?

OpenStudy (anonymous):

-12 ?

Nnesha (nnesha):

yes right

OpenStudy (anonymous):

so there is y = x/-12

OpenStudy (anonymous):

or y = x/12 ?